OpenStudy (michaelgramlick100):

At the bank, Sheila made 6 deposits, each in the same amount. Her sister Sherri made 5 deposits, each in the same amount. Each of Sherri's deposits was $10 more than each deposit Sheila made. Both sisters deposited the same amount in the end. How much did each sister deposit each time? (a) Write an equation. Let x represent the amount of one of Sheila’s deposits. (b) Solve the equation. Show your work. (c) Check your solution. Show your work. (d) State the solution in complete sentences.

OpenStudy (sapphiremoon):

It gave you your workflow right there. Let's write the equations like in step (a). Sheila deposits $x each time. Sherri deposits $10 more, or $x+10 each time. After 6 of Sheila's and 5 of Sherri's deposits, they have the same amount, so: \[6x = 5(x + 10)\] Now, we need to solve the equation as in (b). First we distribute the 5: \[6x = 5x + 50\ Subtract 5x from both sides to get x on only the LHS: \[6x - 5x = 5x -5x + 50\] \[x = 50\] Magic. Now, to check the solution as in (c), just plug your x value into the equation and see if it's true: \[6(50) = 5(50 + 10)\] \[6(50) = 5(60)\] \[300 = 300\] Yay, we got it right! The solution in complete sentences as in (d) is: Sheila deposited $50, and Sherri deposited $60 each time.

OpenStudy (sapphiremoon):

Also, this goes in the Mathematics group, not the Chemistry group.

OpenStudy (michaelgramlick100):

ok thanks i thought i was in the math group

OpenStudy (sapphiremoon):

yw, it's fine, I'm over there, too, anyway.